OpenStudy (calculusxy):

Ms. Diaz wants to divide her class of 30 students into 10 groups. The groups do not need to be of equal size. What are some of her choices? What does "10 groups of 3, 9 of 1, one of 21, there are a lot since they dont have to b even."

OpenStudy (calculusxy):

@JuanitaM

OpenStudy (anonymous):

Basically, you are just finding ways to add up to 30 based on 10 groups..the examples you gave are great!

OpenStudy (calculusxy):

So you have to find the addends that can add up to 30? @JuanitaM

OpenStudy (anonymous):

yes... for instance - you can have 10 groups of 3 people or you can have 1 group of 21 and 9 groups of 1 person each see it - they all total up to 30 groups and you can only account for 30 people

OpenStudy (calculusxy):

How do you get 1 group of 21? @JuanitaM

OpenStudy (anonymous):

putting 21 students in 1 group and then 9 groups of 1 person each...totals 10 groups right?

OpenStudy (calculusxy):

lets say for instance, that I put in 1 group 12 and in the other 9 group I put 2 people each. Is this correct? How did you get 21 and 9? Are they together? @JuanitaM

OpenStudy (anonymous):

ok...1 + 9 = 10 correct 12 + 18 = 30 correct yes, your example works

OpenStudy (anonymous):

yes 21 and 9 are together 1 group of 21 kids 9 groups of 1 person 10 group and 21+9 = 30 kids

OpenStudy (calculusxy):

Thank you for your time! I finally understood!

OpenStudy (anonymous):

See you can do it!!!! Happy studies...